Evaluate the integral
Intergral xsquared +1over square root of xcubed +3x whole thing to dx
using appropriate substitution.
1
Expert's answer
2020-06-14T17:11:11-0400

We can evaluate this integral by substitution.

Set "u=x^{3}+3x". This means "\frac {du}{dx}=3x^{2}+3", or "dx=\frac{du}{3(x^2+1)}."

Now

"\int \frac{x^2+1}{\sqrt{x^3+3x}}dx=\int \frac{x^2+1}{\sqrt{u}} \frac{du}{3(x^2+1)}=
\int \frac{du}{3 \sqrt{u}}="

"=\frac{1}{3} \int \frac{du}{\sqrt{u}}= \frac{1}{3} \cdot 2\sqrt{u}+C=
\frac{2}{3}\sqrt{u}+C="

"=\frac{2}{3}\sqrt{x^3+3x}+C."